OpenStudy (anonymous):

find the line that passes through the point ( 2, 0, 1) and is perpendicular to plane 2x -4y-5z + 8 =0

OpenStudy (amistre64):

strip the normal and apply it to the stated point

OpenStudy (amistre64):

the normal IS perp to the plane .. therefore ....

OpenStudy (anonymous):

the normal is (2 , -4 , -5 ) ? how do I find the line ?

OpenStudy (amistre64):

add some multiple, t, of the normal vector to the state point

OpenStudy (amistre64):

 L = P + n

OpenStudy (amistre64):

the parametric line would be to define xyz in their own right

OpenStudy (amistre64):

x = px + (nx)t y = py + (ny)t z = pz + (nz)t

OpenStudy (amistre64):

this can also be expressed as the matrix product: \[ \begin{pmatrix} px&nx\\ py&ny\\ pz&nz\\ \end{pmatrix} \begin{pmatrix} 1\\t \end{pmatrix} = \begin{pmatrix} x\\y\\z \end{pmatrix} \]

OpenStudy (anonymous):

the nx is the normal vectors ? what is px then?

OpenStudy (amistre64):

point (x,y,z) px is my own shorthand for the x component of the given point ... nx has a similar concept

OpenStudy (anonymous):

Sorry, ok px is the given point? (2,0,1)

OpenStudy (amistre64):

the given point is:( 2, 0, 1) px = 2 py = 0 pz = 1

OpenStudy (anonymous):

Eq. of any line through (x1,y1,z1) with direction ratios a,b,c is \[\frac{ x-x1 }{a }=\frac{ y-y1 }{ b }=\frac{ z-z1 }{c }\] Here (x1,y1,z1)=(2,0,1) a,b,c are proportional to 2,-4,-5 respectively.